# The second class particle in the half-line open TASEP

### v1

@@ 1.2. Multi-color half-line open TASEP with one shock or two shocks
The first half-line open TASEP has three colors (we denote the first class particle by 1, second class particle by 2 and hole by +$\infty$+ ∞), which is defined on the state-space parametrized by occupation variables $\{(\eta(x))_{x\in\mathbb{Z}_{>0}}\in\{1,2,+\infty\}^{\mathbb{Z}_{>0}}\}$. The initial configuration $\eta_{0}(x)$ has one second class particle which stands between $M_{1}$ holes (on the interval $[1,M_{1}]$) and $M_{2}$ first class particles (on the interval $[M_{1}+2,M_{1}+M_{2}+1]$), the positions larger than $M_{1}+M_{2}+1$ are all filled with holes. See Figure 1 upper. The state $\eta_{t}(x)$ at given time $t\in[0,\infty)$ and position $x\in\mathbb{Z}_{>0}$ evolves according to the following dynamics: a particle $\eta_{t}(x)$ jumps from site $x$ to $x+1$ at exponental rate 1 if $\eta_{t}(x)<\eta_{t}(x+1)$, a first class particle is created at site 1 at rate $\alpha$ if $\eta_{t}(1)>1$. Otherwise, we keep the state unchanged. See Figure 2.
The second half-line open TASEP (we denote by $\tilde{\eta}_{t}$) is similar with the first one, but with a different initial condition. The second initial configuration(we denote by $\tilde{\eta}_{0}$) can be obtained from the first one by replacing the first class particles on the interval $[M_{1}+2,M_{1}+M_{2}+1]$ with the second class particles. See Figure 1 bottom. Obviously, the initial conditions $\eta_{0}$ and $\tilde{\eta}_{0}$ create one shock.
@@
We extend the one-shock initial configuration to the two-shock case by the following way. In the two-shock case of the first initial configuration $\eta_{0}$ (we denote by $\xi_{0}$), we put $N$ holes on the interval $[1,N]$, $M$ first class particles on the interval $[N+1,N+M]$, $M$ holes on the interval $[N+M+1,N+2M]$, one second class particle at the position $N+2M+1$, $N$ first class particles on the interval $[N+2M+2,2N+2M+1]$, and the positions larger that $22N+2M+$ 2 are all filled with holes. We denote the half-line open TASEP with initial condition $\xi_{0}$ by $\xi_{t}$. See Figure 3 Upper. The dynamics of $\xi_{t}$ are also given by Figure 2.
The two-shock case of the second initial configuration $\tilde{\eta}_{0}$ (we denote by $\tilde{\xi}_{0}$) can be obtained from $\xi_{0}$ by replacing the particles on the interval $[N+1,N+M]$ with the second class particles, the particles on the interval $[N+2M+1,2N+2M+1]$ with the third class particles (we denote by 3). We denote the half-line open TASEP with initial condition $\tilde{\xi}_{0}$ by $\tilde{\xi}_{t}$. See Figure 3 Bottom. The dynamics of $\tilde{\xi}_{t}$ are also given by Figure 2, but with $k\in\{2,3,+\infty\}$, $i<j$ and $i,j\in\{1,2,3,+\infty\}$.
@@ 1.3. Asymptotic distribution of the second class particles under the constant scaling
For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we define the DEHP-tree of $\eta$ by the following way:
@@
Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
@@
The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
(1.1) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n},$
when $\alpha\leq\frac{1}{2}$, where we use the binomial coefficient $\binom{p}{k}=\frac{p!}{k!(p-k)!}$; and it is given by
(1.2) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$.
@@
(1.3) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(1.4) $\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ have the following relationships:
(1.5) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=M_{2}+1).$
The asymptotic distribution of the unique second class particle disappearing from $\eta_{t}$ is given by the following formula:
(1.6) $\lim_{t\to\infty}\mathbb{P}(f(t)<0)=\sum_{m=0}^{M_{2}}\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m).$
Theorem 1.3 (Theorem 4.10).
The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
(1.7) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(1.8) $\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$, where we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$.
@@
(1.9) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}$
when $\alpha\leq\frac{1}{2}$, and it is given by
(1.10) $\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)$
when $\alpha>\frac{1}{2}$.
The asymptotic distribution of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ have the following relationships:
(1.11) $\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=N+1).$
The asymptotic distribution of the unique second class particle disappearing from $\xi_{t}$ is given by the following formula:
(1.12) $\lim_{t\to\infty}\mathbb{P}(g(t)<0)=\sum_{s=0}^{N}\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s).$
@@
The asymptotic distribution of the height function of second class particles in $\tilde{\xi}_{t}$ is similar with (1.3) and (1.4). For the sake of brevity, we only provide the asymptotic distribution of the height function of third class particles in $\tilde{\xi}_{t}$. The third class particles can be regarded as the second-type second class particles.
@@ 1.4. Asymptotic distribution of the second class particles under the KPZ scaling
In this section, we consider the asymptotic probability distribution of the second class particles under the KPZ scaling $M_{1}=M_{2}=\lfloor at^{\frac{2}{3}}\rfloor$, $a\in\mathbb{R}_{>0}$. Below we will use the half-space ${Airy}_{2}$ process $\mathcal{A}_{\varpi}^{HS}$ which was introduced in [BBCS18a], see section 5 for the definition. Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
@@ 2.1. Colored half-space TASEP
The dynamics of this model involve associating exponential clocks of rate 1 with each edge $(x,x+1)$ for $x\geq 1$, and an exponential clock of rate $\alpha$ with the edge $(-1,1)$. When a clock attached to the edge $(x,x+1)$ with rate 1 rings, particles are swapped at positions $x$ and $x+1$ if $\eta(x)<\eta(x+1)$, and a similar swap occurs for particles at positions $-x$ and $-x-1$ at the same time. When the clock with rate $\alpha$ rings, a swap occurs between $\eta(-1)$ and $\eta(1)$ if $\eta(-1)<\eta(1)$. Otherwise, we remain the configuration of system unchanged. See Figure 5 for an illustration. For any fixed time $t$, there is almost surely in finding a sequence $i_{k}\in\mathbb{N}$ such that no clock associated with the edge $(i_{k},i_{k}+1)$ rings up to time $t$. Given this sequence, the process can be sampled within the finite intervals $[i_{k}+1,i_{k+1}]$, ensuring the well-defined nature of this process (see [Har78]).
@@ 2.2. Random walk on the type BC Hecke algebra
which can be straightforwardly proved by induciton in $l(\omega)$ with the use of (2.1). 2.2 was first introduced in probabilistic setting by [Buf20]. Furthermore, we use the fact for any Coxeter generator $\iota(T_{s_{k}})=T_{s_{k}}$ to induce the following lemma:
@@ 2.3. Color-position symmetry theorem
$\iota(\eta_{t}^{gen;2})=\iota(T_{s_{k_{1}}}\cdots T_{s_{k_{n}}}W(t)T_{id}),$

### v2

@@ 1.2. Multi-color half-line open TASEP with one shock or two shocks
The first half-line open TASEP has three colors (we denote the first class particle by 1, second class particle by 2 and hole by +$\infty$+ ∞), which is defined on the state-space parametrized by occupation variables $\{(\eta(x))_{x\in\mathbb{Z}_{>0}}\in\{1,2,+\infty\}^{\mathbb{Z}_{>0}}\}$. The initial configuration $\eta_{0}(x)$ has one second class particle which stands between $M_{1}$ holes (on the interval $[1,M_{1}]$) and $M_{2}$ first class particles (on the interval $[M_{1}+2,M_{1}+M_{2}+1]$), the positions larger than $M_{1}+M_{2}+1$ are all filled with holes. See Figure 1 upper. The state $\eta_{t}(x)$ at given time $t\in[0,\infty)$ and position $x\in\mathbb{Z}_{>0}$ evolves according to the following dynamics: a particle $\eta_{t}(x)$ jumps from site $x$ to $x+1$ at exponental rate 1 if $\eta_{t}(x)<\eta_{t}(x+1)$, a first class particle is created at site 1 at rate $\alpha$ if $\eta_{t}(1)>1$. Otherwise, we keep the state unchanged. See Figure 2. Let $f(t)$ represent the position of the unique second class particle in the process $\eta_{t}$.
The second half-line open TASEP (we denote by $\tilde{\eta}_{t}$) is similar with the first one, but with a different initial condition. The second initial configuration(we denote by $\tilde{\eta}_{0}$) can be obtained from the first one by replacing the first class particles on the interval $[M_{1}+2,M_{1}+M_{2}+1]$ with the second class particles. See Figure 1 bottom. Obviously, the initial conditions $\eta_{0}$ and $\tilde{\eta}_{0}$ create one shock. Let $\mathbf{N}_{c}(x,t)$ denote the number of particles of color $c$ in the process $\tilde{\eta}_{t}$ at time $t$ that are located weakly to the right of $x\in\mathbb{R}_{\geq 0}$, where $c=1,2$.
@@
We extend the one-shock initial configuration to the two-shock case by the following way. In the two-shock case of the first initial configuration $\eta_{0}$ (we denote by $\xi_{0}$), we put $N$ holes on the interval $[1,N]$, $M$ first class particles on the interval $[N+1,N+M]$, $M$ holes on the interval $[N+M+1,N+2M]$, one second class particle at the position $N+2M+1$, $N$ first class particles on the interval $[N+2M+2,2N+2M+1]$, and the positions larger that $22N+2M+$ 2 are all filled with holes. We denote the half-line open TASEP with initial condition $\xi_{0}$ by $\xi_{t}$. See Figure 3 Upper. The dynamics of $\xi_{t}$ are also given by Figure 2. We denote the position of the unique second class particle in the process $\xi_{t}$ by $g(t)$.
The two-shock case of the second initial configuration $\tilde{\eta}_{0}$ (we denote by $\tilde{\xi}_{0}$) can be obtained from $\xi_{0}$ by replacing the particles on the interval $[N+1,N+M]$ with the second class particles, the particles on the interval $[N+2M+1,2N+2M+1]$ with the third class particles (we denote by 3). We denote the half-line open TASEP with initial condition $\tilde{\xi}_{0}$ by $\tilde{\xi}_{t}$. See Figure 3 Bottom. The dynamics of $\tilde{\xi}_{t}$ are also given by Figure 2, but with $k\in\{2,3,+\infty\}$, $i<j$ and $i,j\in\{1,2,3,+\infty\}$. Let $\mathsf{N}_{i}(x,t)$ be the number of particles of color $i$ in $\tilde{\xi}_{t}$ which are weakly to the right of $x\in\mathbb{R}$, where $i\in\{1,2,3\}$.
@@ 1.3. Asymptotic distribution of the second class particles under the constant scaling
Let $\alpha\in(0,1)$ and $c=\frac{1}{4}$. For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we define the DEHP-tree of $\eta$ by the following way:
@@
Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha\leq\frac{1}{2}$:
@@
Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
$\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\binom{M_{1}+M_{2}+1}{n}\alpha^{n}(1-\alpha)^{M_{1}+M_{2}+1-n}.$
@@
$\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\binom{M_{1}+M_{2}+1}{M_{1}+M_{2}+1-m}\alpha^{M_{1}+M_{2}+1-m}(1-\alpha)^{m}.$
Our second main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$ when $\alpha>\frac{1}{2}$:
Theorem 1.3 (Theorem 4.11 modulo Conjecture 4.3).
Let $\alpha>\frac{1}{2}$ and we use the notation $B_{n}=\{\eta\in\{0,1\}^{M_{1}+M_{2}+1}:\ell(\eta)=n\}$. The asymptotic distribution of the unique second class particle existing in $\eta_{t}$ is given by:
$\lim_{t\to\infty}\mathbb{P}(f(t)>0)=\sum_{n=0}^{M_{1}}\sum_{\eta\in B_{n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
$\lim_{t\to\infty}\mathbb{P}(\mathbf{N}_{2}(1,t)=m)=\sum_{\eta\in B_{M_{1}+M_{2}+1-m}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
Our third main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha\leq\frac{1}{2}$:
Theorem 1.4 (4.12).
Let $\alpha\leq\frac{1}{2}$. The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
$\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\binom{M+N}{m}\binom{M+N+1}{n}\alpha^{m+n}(1-\alpha)^{2M+2N+1-m-n}.$
@@
$\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\binom{M+N}{N-n}\binom{M+N+1}{M+N+1-s+n}\alpha^{M+2N+1-s}(1-\alpha)^{M+s}\\ +\sum_{m=N}^{M+N}\binom{M+N}{m}\binom{M+N+1}{M+N+1-s}\alpha^{M+N+1+m-s}(1-\alpha)^{M+N-m+s}.$
Our fourth main result is the asymptotic behavior of the second class particle in $\xi_{t}$ and the third class particles in $\tilde{\xi}_{t}$ when $\alpha>\frac{1}{2}$:
Theorem 1.5 (Theorem 4.13 modulo Conjecture 4.3).
Let $\alpha>\frac{1}{2}$ and we use the notation $B_{m,n}=\{\eta\in\{0,1\}^{2M+2N+1}:\sum_{i=1}^{M+N}\eta_{i}=m,\sum_{j=M+N+1}^{2M+2N+1}\eta_{j}=n\}$. The asymptotic distribution of the unique second class particle existing in the half-space TASEP $\xi_{t}$ is given by:
$\lim_{t\to\infty}\mathbb{P}(g(t)>0)=\sum_{m=0}^{N-1}\sum_{n=0}^{M+N-m}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{n=0}^{M}\sum_{\eta\in B_{m,n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
$\lim_{t\to\infty}\mathbb{P}(\mathsf{N}_{3}(1,t)=s)=\sum_{n=1}^{s}\sum_{\eta\in B_{N-n,M+N+1-s+n}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right)\\ +\sum_{m=N}^{M+N}\sum_{\eta\in B_{m,M+N+1-s}}\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
@@
The asymptotic distribution of the height function of second class particles in $\tilde{\xi}_{t}$ is similar with (4.13) and (4.15). For the sake of brevity, we only provide the asymptotic distribution of the height function of second class particles in $\tilde{\xi}_{t}$.
@@ 1.4. Asymptotic distribution of the second class particles under the KPZ scaling
In this section, we consider the asymptotic probability distribution of the second class particles under the KPZ scaling $t^{\frac{2}{3}}$. Below we will use the half-space ${Airy}_{2}$ process $\mathcal{A}_{\varpi}^{HS}$ which was introduced in [BBCS18a], see section 5 for the definition. Our first main result is the asymptotic behavior of the second class particles in $\eta_{t}$ and $\tilde{\eta}_{t}$:
@@ 2.1. Colored half-space TASEP
The dynamics of this model involve associating exponential clocks of rate 1 with each edge $(x,x+1)$ for $x\geq 1$, and an exponential clock of rate $\alpha$ with the edge $(-1,1)$. When a clock attached to the edge $(x,x+1)$ with rate 1 rings, particles are swapped at positions $x$ and $x+1$ if $\eta(x)<\eta(x+1)$, and a similar swap occurs for particles at positions $-x$ and $-x-1$ at the same time. When the clock with rate $\alpha$ rings, a swap occurs between $\eta(-1)$ and $\eta(1)$ if $\eta(-1)<\eta(1)$. Otherwise, we leave the system configuration unchanged. See Figure 5 for an illustration. For any fixed time $t$, there is almost surely in finding a sequence $i_{k}\in\mathbb{N}$ such that no clock associated with the edge $(i_{k},i_{k}+1)$ rings up to time $t$. Given this sequence, the process can be sampled within the finite intervals $[i_{k}+1,i_{k+1}]$, ensuring the well-defined nature of this process (see [Har78]).
@@ 2.2. Random walk on the type BC Hecke algebra
Remark 2.1.
While the constructions in this section describe stochastic processes on a finite space, they can be extended to the infinite setting using the standard argument of Harris ([Har78]). Namely, at any fixed time, the processes on an infinite space are, with probability 1, (infinite) collections of finite-space processes that did not interact with each other. Consequently, the results for processes on an infinite space follow directly from the corresponding results for their finite-space counterparts.
Remark 2.2.
Since the random element of the Hecke algebra and the corresponding random permutation that it induces are equivalent, for convenience, we will use the two concepts equivalently in the following descriptions and do not mention it again.
@@
which can be straightforwardly proved by induction in $l(\omega)$ with the use of (2.1). 2.2 was first introduced in probabilistic setting by [Buf20]. Furthermore, we use the fact for any Coxeter generator $\iota(T_{s_{k}})=T_{s_{k}}$ to induce the following lemma:
@@ 2.3. Color-position symmetry theorem
$\iota(\eta_{t}^{gen;2})=\iota(T_{s_{k_{1}}}\cdots T_{s_{k_{n}}}W(t)),$